Your nearsighted instructor cannot focus clearly on objects more distant than 150 cm from her eyes. What power lenses are required for her to see distant students clearly?
Answer #1

Given :

Far point is 150 cm.

So to see distant students clearly, the lens to be used forms a virtual image of the distance object at 150 cm so, that she can see distant students.

so, u = object distance = infinity

v = image distance = -150 cm

f = focal length of lens

P = power of lens = \frac{1}{f} [f in meter]

Using lens equation :

\frac{1}{f}=\frac{1}{v}-\frac{1}{u}

\Rightarrow \frac{1}{f}=\frac{1}{-150}-\frac{1}{-\infty} = -\frac{1}{150}

=> f = -150 cm = -1.5 m

therefore, Power of lens to be used = P = -1/1.5 = -0.67D [answer]